OpenStudy (anonymous):

What is the average binding energy per nucleon for a C-12 nucleus with a mass defect of 0.0993 amu? (1 amu= 1.66 x 10-27 kg; 1 J = 1 kg m2/s2)

OpenStudy (anonymous):

I understand it uses E=MC^2 somehow, but I'm not quite sure the exact way it does.

OpenStudy (unklerhaukus):

convert the mass defect to kilograms \[D=0.0993 [\text{amu}]\\\qquad\qquad=0.0993 [\text{amu}]\times1.66 \times 10^{-27} \frac{[\text{kg}]}{[\text{amu}]}\\\qquad\qquad\qquad=\dots[\text{kg}]\] then convert to energy \[\quad=\dots[\text{kg}]\times{\big(c[\text{m/s}]\big)^2}\\ \,\\ \qquad=\dots\times{c^2}[\text{J}]\] If you want energy in terms of electron volts \[\qquad=\dots\times{c^2} [\text{J}]\div{1.60\times10^{-19}}\frac{[\text{J}]}{[\text{eV}]}\\ \qquad=\dots\times{c^2}\div{1.60\times10^{-19}}{[\text{eV}]}\\ \qquad=\] Divide this by the number of nucleons to find the average binding energy per nucleon